# Stein Boltzmann Sampling: A Variational Approach for Global Optimization

### 20 Feb 2024 v2

@@ Abstract
In this paper, we introduce a new flow-based method for global optimization of Lipschitz functions, called Stein Boltzmann Sampling (sbs). Our method samples from the Boltzmann distribution that becomes asymptotically uniform over the set of the minimizers of the function to be optimized. Candidate solutions are sampled via the Stein Variational Gradient Descent algorithm. We prove the asymptotic convergence of our method, introduce two sbs variants, and provide a detailed comparison with several state-of-the-art global optimization algorithms on various benchmark functions. The design of our method, the theoretical results, and our experiments, suggest that sbs is particularly well-suited to be used as a continuation of efficient global optimization methods as it can produce better solutions while making a good use of the budget.
Global optimization, SVGD, Boltzmann distribution, Stein discrepancy, Lipschitz functions, Asymptotic uniform distribution, Consistency, Particle filtering, Simulated Annealing
@@ 1 Introduction
Sequential methods rely on two components: a sampling process to explore the search space, and a selection process to choose the next candidate solution using the information given by the previous samples. In this work, we introduce a new sequential, flow-based and deterministic method called Stein Boltzmann Sampling (sbs) for Lipschitz functions. Our method uses the Stein Variational Gradient Descent (svgd) (Liu & Wang, 2016) method to sample from the Boltzmann distribution, which has the characteristic that tends to a uniform distribution over the set of the minimizers. svgd constructs a flow in the space of probability measures (similarly to the way a gradient flow would evolve in $\mathbb{R}^{d}$) that moves towards the target sampling measure. Even though our method is not a typical stochastic one (since svgd sampling is deterministic), we prove its asymptotic convergence for any Lipschitz function using elements of the svgd theory. We show that the sbs method achieves competitive performance on standard global optimization benchmarks versus three stochastic state-of-the-art methods. The first one, adalipo (Malherbe & Vayatis, 2017), is consistent over Lipschitz functions and is adapted for a very low computational budget (i.e. function evaluations at candidate minimizers). The second and third ones, cma-es (Hansen & Ostermeier, 1996, 2001; Hansen et al., 2003) and woa (Mirjalili & Lewis, 2016), are two inconsistent methods but known to be very efficient in practice. Due to either early stopping conditions or time complexity, these three existing methods do not scale computationally well, hence they are not suited for when the available computational budget is low.
@@ 2.1 The proposed method
We introduce the Stein Boltzmann Sampling (sbs) method. The idea is to sample from a distribution that converges asymptotically to a uniform distribution over the set of minimizers $X^{*}$ of an arbitrary continuous function $f$. We use the continuous Boltzmann distribution (BD) for this purpose.
@@
A characteristic property of the BD is that it tends to a uniform distribution over the set of minimizers $X^{*}$ as $\kappa$ tends to infinity (see Figure 1). More details can be found in Section 3.1. The sbs method aims to sample from the BD with $\kappa$ large enough in order for the function values at the sampled points to be close to the global minimum. As $m^{(\kappa)}$ converges to a uniform distribution over $X^{*}$, the approximation of $f^{*}$ can be made arbitrarily accurate. However, as it is not efficient to sample from BD by estimating the intractable term $\int_{\Omega}e^{-\kappa f(t)}\mathrm{d}t$ using classical Monte-Carlo methods, we propose to use instead the Stein Variational Gradient Descent (svgd) method. Given an initial measure $\mu$, svgd constructs iteratively a flow of measures that moves towards the target measure, noted as $\pi$. The update direction is given by:
@@ 2.3 Discrete setting
Let $f:\Omega\to\mathbb{R}$ be a Lipschitz function. Let $\kappa>0$ and let $\pi$ be the BD defined in Definition 2.1 associated with $f$ and $\kappa$. Let $\mu\in\mathcal{P}_{2}(\Omega)$ such that $\mu\ll\lambda$, $\mathrm{supp}(\mu)=\Omega$ and $K\!L(\mu||\pi)<\infty$. Let $\hat{\mu}_{n}$ be defined by Equation 3. Then, when $\varepsilon\to 0$,
$\left\{f\left(X^{(n)}\right)\;\middle|\;X^{(n)}=(x_{1},\dots,x_{N})\sim\hat{\mu}^{\otimes N}_{n}\right\}\xrightarrow[\begin{subarray}{c}\kappa\to\infty\\ N\to\infty\\ n\to\infty\end{subarray}]{\mathcal{L}}\{f^{*}\}.$
@@ 3.1 Boltzmann distribution
Recall that the BD has been formally defined in Definition 2.1. The BD is a well-known distribution in statistical physics. It is used to model the distribution of the energy of a system in thermal equilibrium. The parameter $\kappa$ is called the inverse temperature. The higher $\kappa$ is, the more concentrated the mass is around the minima of $f$. When $\kappa$ tends to infinity, the BD tends to a uniform distribution over the minima of $f$. The BD is typically used in a discrete settings, i.e. where the number of states is finite. The continuous version can be defined using the Gibbs measure. The following properties come from (Luo, 2019). For the sake of completeness, we provide the proofs in Section A.1.

### 28 Feb 2024 v3

@@ Abstract
In this paper, we introduce a new flow-based method for global optimization of Lipschitz functions, called Stein Boltzmann Sampling (sbs). Our method samples from the Boltzmann distribution that becomes asymptotically supported over the set of the minimizers of the function to be optimized. Candidate solutions are sampled via the Stein Variational Gradient Descent algorithm. We prove the asymptotic convergence of our method, introduce two sbs variants, and provide a detailed comparison with several state-of-the-art global optimization algorithms on various benchmark functions. The design of our method, the theoretical results, and our experiments, suggest that sbs is particularly well-suited to be used as a continuation of efficient global optimization methods as it can produce better solutions while making a good use of the budget.
Global optimization, SVGD, Boltzmann distribution, Stein discrepancy, Lipschitz functions, Asymptotic distribution, Consistency, Particle filtering, Simulated Annealing
@@ 1 Introduction
Sequential methods rely on two components: a sampling process to explore the search space, and a selection process to choose the next candidate solution using the information given by the previous samples. In this work, we introduce a new sequential, flow-based and deterministic method called Stein Boltzmann Sampling (sbs) for Lipschitz functions. Our method uses the Stein Variational Gradient Descent (svgd) (Liu & Wang, 2016) method to sample from the Boltzmann distribution, which has the characteristic that tends to a distribution supported over the set of the minimizers. svgd constructs a flow in the space of probability measures (similarly to the way a gradient flow would evolve in $\mathbb{R}^{d}$) that moves towards the target sampling measure. Even though our method is not a typical stochastic one (since svgd sampling is deterministic), we prove its asymptotic convergence for any Lipschitz function using elements of the svgd theory. We show that the sbs method achieves competitive performance on standard global optimization benchmarks versus three stochastic state-of-the-art methods. The first one, adalipo (Malherbe & Vayatis, 2017), is consistent over Lipschitz functions and is adapted for a very low computational budget (i.e. function evaluations at candidate minimizers). The second and third ones, cma-es (Hansen & Ostermeier, 1996, 2001; Hansen et al., 2003) and woa (Mirjalili & Lewis, 2016), are two inconsistent methods but known to be very efficient in practice. Due to either early stopping conditions or time complexity, these three existing methods do not scale computationally well, hence they are not suited for when the available computational budget is low.
@@ 2.1 The proposed method
We introduce the Stein Boltzmann Sampling (sbs) method. The idea is to sample from a distribution that converges asymptotically to a distribution supported over the set of minimizers $X^{*}$ of an arbitrary continuous function $f$. We use the continuous Boltzmann distribution (BD) for this purpose.
@@
A characteristic property of the BD is that it tends to distribution supported over the set of minimizers $X^{*}$ as $\kappa$ tends to infinity. If $\lambda(X^{*})>0$, the BD tends to a uniform distribution over $X^{*}$ (see Figure 1). If $X^{*}$ is finite, it tends to a sum of Dirac distribution over $X^{*}$ where the weight on each minimizer depends on the local geometry of the function (Hwang, 1980). More details can be found in Section 3.1. The sbs method aims to sample from the BD with $\kappa$ large enough in order for the function values at the sampled points to be close to the global minimum. As $m^{(\kappa)}$ converges to a distribution supported over $X^{*}$, the approximation of $f^{*}$ can be made arbitrarily accurate. However, as it is not efficient to sample from BD by estimating the intractable term $\int_{\Omega}e^{-\kappa f(t)}\mathrm{d}t$ using classical Monte-Carlo methods, we propose to use instead the Stein Variational Gradient Descent (svgd) method. Given an initial measure $\mu$, svgd constructs iteratively a flow of measures that moves towards the target measure, noted as $\pi$. The update direction is given by:
@@ 2.3 Discrete setting
Let $f:\Omega\to\mathbb{R}$ be a Lipschitz function. Let $\kappa>0$ and let $\pi$ be the BD defined in Definition 2.1 associated with $f$ and $\kappa$. Let $\hat{\mu}_{0}\in\mathcal{P}_{2}(\Omega)$ such that $\hat{\mu}_{0}\ll\lambda$, $\mathrm{supp}(\hat{\mu}_{0})=\Omega$ and $K\!L(\hat{\mu}_{0}||\pi)<\infty$. Let $\hat{\mu}_{n}$ be defined by Equation 3. Then, when $\varepsilon\to 0$,
$\left\{f\left(X^{(n)}\right)\;\middle|\;X^{(n)}=(x^{(1)},\dots,x^{(N)})\sim\hat{\mu}^{\otimes N}_{n}\right\}\xrightarrow[\begin{subarray}{c}\kappa\to\infty\\ N\to\infty\\ n\to\infty\end{subarray}]{\mathcal{L}}\{f^{*}\}.$
@@ 3.1 Boltzmann distribution
Recall that the BD has been formally defined in Definition 2.1. The BD is a well-known distribution in statistical physics. It is used to model the distribution of the energy of a system in thermal equilibrium. The parameter $\kappa$ is called the inverse temperature. The higher $\kappa$ is, the more concentrated the mass is around the minima of $f$. When $\kappa$ tends to infinity, the BD tends to a distribution supported over the minima of $f$. The BD is typically used in a discrete settings, i.e. where the number of states is finite. The continuous version can be defined using the Gibbs measure. The following properties come from (Luo, 2019). For the sake of completeness, we provide the proofs in Section A.1.